OpenStudy (anonymous):

Write 19 and 1/2 in radical form

jimthompson5910 (jim_thompson5910):

you mean \[\Large 19^{1/2}\] right?

jimthompson5910 (jim_thompson5910):

the 1/2 is an exponent

OpenStudy (anonymous):

Yes what you said jim

jimthompson5910 (jim_thompson5910):

so it's simply \[\Large 19^{1/2} = \sqrt[2]{19} = \sqrt{19}\] the index of '2' is not needed since square roots are by default radicals with index of 2

jimthompson5910 (jim_thompson5910):

I'm using the rule \[\LARGE x^{1/n} = \sqrt[n]{x}\]